Prev Next

SW&IT Meeting Minutes 2019-05-27

Attendees: Memorial Day

Outlook

Labor Reporting

Report auto-generated every, Do not edit this table, use entries in smartsheets.

News from management

  • yearly evaluations

Previous Action Items

  • JH "Not getting all ITs"
  • lab jumbo frame (IT)
  • announcing and delivering of new s/w

Open topics

  • wiki need for an admin/mod

Team reports

MB

  • OC, WPC, Management
  • new dss rpm
  • f30 ci and rpm build catch-up
  • dms telemetry, x/y, psd plots
  • looking at 128Hz vibration with dms (IT)
  • discuss misc s/w with JHill
  • resucitate log.io
next
  • obs1/f30 catch-up loose ends
  • f30 godzilux w/s delivery follow up
  • check freshdesk for MWagner
  • AO priorities

PK

  • Ice3.7
  • gdb "catch throw"
  • Readline history - *using_history*, *add_history*, *write_history*, *read_history*
  • TCS 2019A testing - problems with MODS guiding probes, general problems with rotator angles (either caused by new Poitning Kernel, or by my changes..TBD)
  • diff -w (see git help diff)

AC

  • Tried to build tcsweb on CentOS 6 CI and rm507-nuc but without success. To be continued.
  • Built OVMS-INDI
  • Refreshed self with python pandas to bin timeseries data
  • "Smarter" FACSUM wip - show status that the network is down

YZ

  • LBC flat wiki updates:
    • Completed an initial Python 3 program that does everything (scanning LBC images to find new flats, as well as formatting and updating wiki page).
    • Currently set it up on the 'ao-cravat' VM as a cron task (running once a day). Olga informed me that there may be new LBC flats coming on May 29/30. I will monitor its performance.
    • Decided to put it under github repo: swg-tools/lbcflatwiki. Created a pull request to merge it.
    • Some further work planned include:
      • Improve its efficiency and ensure correctness.
      • Refactor the code, especially maybe separate the wiki login and updates into a new Python module.
  • TCS:
    • Petr helped to setup a new MCSPU simulator.
    • Xianyu helped to test new SHARK authorization with the real AO supervisor VM.
    • We have concluded that the new SHARK authorization should not have any problems. However we saw some other errors during the simulation we didn't fully understand yet. Plan to investigate them a bit further when I have some time.
    • Next step is to create two pull requests:
      • For SHARK authorization
      • For some updates to the ApplyWaffleAO commands (mainly small corrections and setting up IIF event handles).
  • Some more online safty trainings.

DE

  • Web work
  • 4D support
    • Tucson IPs
    • Light Ethernet ports
  • UITS colo quote refresh
  • Misc. office purchases
  • Got Louise to move phone list to GDocs, changed wiki link
  • John Hill IT emails blocked(?) by astronomy department mail server? LBTO email works, I have changed his IT user account to use that one for now.

SH

  • Obs1 upgraded -- good for Josh, pending Steve A. (talk to him)
  • WiFi on level 4. Lit and tested.
  • Measured R3.
  • Played with mountain emergency response debugging entity.
  • ICE 3.7 installed on tcs{1..3} and tcs-test.
next:
  • VM for catalogs.mountain
  • Look at Steve A. and John H. scripts.
  • some IRAF cleanup -- LBTtools.
  • Re-work some of kickstart.

XZ

Last week
  • Starting working on AO SW for 4D
  • SHARK authorization and AO supervisor simulator test with Yang
  • SOUL commnunication issue
  • Working on DX AdSec disturbance cannot apply issue
  • SX AdSec strange status issue
  • OVMS document
This week
  • Contiue working on 4D
  • AdSec issues

IS

Last week
  • LUCI
    • Agreed to partially revert fix for #7642 (RTD aladin coordinate reading on X/Y has 0.5 px mismatch). I was adding a 0.5px for every call to aladin, we are not sure how this will affect the tool, so I'll just leave the original fix which only affects nonsidereal longslits. Need to clean up.
    • Working on issue #7473. Most of the code is done and tested. Need to clean up.
    • Discuss with Dave and Michelle a plan to implement SRR and subframing for LUCI
  • OT
    • Went through all open OT issues with Michelle and Andrew Cardwell and got a plan and a timeline to fix them.

This week
  • LUCI
    • Finish off issues #7642 and #7473
    • Also add a fix on a possible issue I found on the backend AFC code.
    • In the same area see if I have find why the Compensation Mirror backend daemon stalls, but wont spend much time on this
    • Finish the luci GUI run scripts in observer machines. See if they can be run from a read only directory.
  • PIT
    • Make sure all scripts are in git
    • clean up Italian spreadsheet script generation.
Topic revision: r16 - 28 May 2019, IgorSola
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback